OpenStudy (anonymous):

Mr. Mendez had 24 cows on his farm. At an auction, he bought 5 more cows. Find the percent of change in the number of cows on the farm. Round your answer to the nearest tenth, if necessary. a. –20.8% c. 79.2% b. 82.8% d. 20.8%

OpenStudy (amoodarya):

first was 24 cows then 24+5 so 5/24 *100 is percentage=0.2083333 *100=20.83 so round this implicit : it is +20.83 %